'Bhartdarshan' is an Internet-based travel agency wherein customers can see videos of the cities they plan to visit. The number of hits daily is a normally distributed random variable with a mean of 10,000 and a standard deviation of 2,400.


a. What is the probability of getting more than 12,000 hits ?



b. What is the probability of getting fewer than 9,000 hits ?


1
Expert's answer
2022-03-29T12:07:04-0400

"\mu=10000,\sigma=2400"

a)

"P(x>12000)=P(Z>\frac{12000-10000}{2400})"

"=P(Z>0.83)"

"=1-P(Z<0.83)"

"=1-0.79673"

"=0.20327"


b)

"P(X<9000)=P(Z<\frac{9000-10000}{2400})"

"P(Z<-0.42)"

"=0.33724"